OpenStudy (anonymous):

integrate 1/(1+x)^3 dx from x=0 to infinity

OpenStudy (kc_kennylau):

\[\Large\int_0^\infty\frac1{(1+x)^3}dx\]

OpenStudy (kc_kennylau):

What is your problem?

Miracrown (miracrown):

what are your thoughts?

OpenStudy (kc_kennylau):

It is a problem in integrating or problem in substituting?

OpenStudy (anonymous):

solve it step by step and integrate 1/(x*lnx) dx from x=0 to infinity integrate 1/(x+3)^(3/2) dx from x=3 to 4 and find the interval or state they diverge

Miracrown (miracrown):

 if you convert to exponential form, you can use the formula of integration i wrote down.. are you familiar with that conversion?